4 pts Question 2 What is the relative frequency of nurses using at least 4 but no more than 6 sick days? (Enter as a decimal rounded to two decimal places)
Window Help abi-FX @ Course c x|囚Ausif Ma × Expedia × 目Untitled × 01%20-%20Frequency%20Distributions%20&%20Histograms%20 Tables.pdf | expe。 TABLE 2.10 Nurse Sick Nurse Sick Nurse Sick Number Days Numher Days Number Days 10 6 18 12 4 13 7 20 TABLE 2.11 App. App Numb. Gender Ws Wp Numb. Gender Ws Wp m 220 225 f 135 137 f 180201 m 210 205 f 145 144 f 131 133 m 177 180 m 135 135 m 183 180 m 165 166 f 160 166 m 178 180 m 155 152 f 155 154 130 128 m 165 167 m 215 210 17 m 190 186 18 m 158 160 20 165 160 21 f 120 118 22 m 173 170 23 m 188 195 24 m 180 195 25 f 135 137 26 m 155 155 27 m 190 19528 8 m 187 154 156 185 29 9 stats.pdf
Answer #1

Ans:

There are 6 nurses who are using at least 4 but no more than 6 (i.e. 4,5,6 sick days)

(nurse no. 5,6,7,11,19,21)

Total number of nurses=21

Relative frequency=6/21=2/7=0.29